# X-ray free electron laser SPI single-particle diffractive imaging

### v1

@@ 4 Experiments
In the following, we describe the computational experiments we perform to arrive at our results and conclusions. For clarity, we preface this section with the following simplifying assumptions we make for computational viability in our simulated setting:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
Relaxations of these assumptions should be possible within our framework through extensions of our formalism, but are left for future work here due to the difficulty that this simplified setup already exhibits, and due to the likely need for parameter tuning in more complex settings.
@@ 4.1 Problem Setup
4.1.1 Forward Model
We treat here a simplified 2D far-field real-space ptychography problem. Let $x\in\mathbb{C}^{H\times W}$ be an image of the complex object transmission function which we want to recover,
$y_{k}$ $=\left|\mathcal{A}(r_{k},x)\right|^{2}+\varepsilon_{k},$ (4.2)
where $\varepsilon_{k}$ is the measurement noise of the $k$-th measurement, $\mathcal{F}$ models the wavefront propagation to the detector and $\operatorname{DFT}$ denotes the discrete Fourier transform, the probe array $p\in\mathbb{C}^{H_{p}\times W_{p}}$ models the complex-valued wavefield of the probe in the plane of the object and is applied through element-wise multiplication $\odot$, and $S(r_{k},\cdot)$ is a shift operator depending on a two-dimensional shift $r_{k}=(\Delta h_{k},\Delta w_{k})$ which we refer to as the scan position in the following. The Crop operator crops its input to the same array size as the probe $p$ before element-wise multiplication. The Pad operator pads its input with one full probe array size of empty space entries; see Section 4.1.2. In all experiments, we use an object array of size $x\in\mathbb{C}^{H\times W}:=\mathbb{C}^{256\times 256}$ and a probe array of size $p\in\mathbb{C}^{H_{p}\times W_{p}}:=\mathbb{C}^{512\times 512}$.
@@ 4.1.5 Simulated Measurements
where each pixel in $y_{k}$ is treated as both the mean and the variance, and the square root is taken element-wise. An alternative would be to use $\left|\mathcal{A}(r_{k},x)\right|^{2}$ as the variance of the approximating Gaussian instead of $y_{k}$, but in preliminary experiments we found this option to lead to less stable reconstructions and hence always use Eq. 4.8. In practice, within the expression $\frac{1}{2y_{k}}$ we clamp $y_{k}$ to have a minimum value of 1 to avoid both division by zero and an overly high weighting of detector pixels with no measured photons.
@@ 4.2.1 Variational Inference
For our variational inference methods and all choices of prior distribution, we approximate the posterior by the following variational distributions:
$q_{\chi}(x|y):=\mathcal{N}_{\mathbb{C}}(x;\mu_{\chi},\Sigma_{\chi}I),$ $\ \mu_{\chi}\in\mathbb{C}^{d},$ $\ \Sigma_{\chi}=\operatorname{diag}(\sigma_{\chi}^{2}),$ $\ \sigma_{\chi}\in\mathbb{R}_{+}^{d}\,$, (4.9)
$q_{\rho}(r|y):=\mathcal{N}(r;\mu_{\rho},\Sigma_{\rho}I),$ $\ \mu_{\rho}\in\mathbb{R}^{K\times 2},$ $\ \Sigma_{\rho}=\operatorname{diag}(\sigma_{\rho}^{2}),$ $\ \sigma_{\rho}\in\mathbb{R}_{+}^{K}\,$, (4.10)
$q_{\chi,\rho}(x,r|y):=q_{\chi}(x|y)q_{\rho}(r|y)\,.$ (4.11)
@@
For both the image and position parameters, we use the Adam optimizer [42], which has previously been established as a working method for (non-blind) ptychographic image reconstruction [39, 37, 33]. We set its momentum parameters to $\beta_{1}=0.9,\beta_{2}=0.999$, the default settings in the PyTorch package [55]. We run $N=\text{10,000}$ outer steps unless otherwise noted, each with 1 inner image optimization step and 10 inner position optimization steps. For the position parameters we set the step sizes $\lambda_{\mu_{\rho}}=10.0$ and $\lambda_{\sigma_{\rho}}=0.01$, where we use a high learning rate of the means ($\lambda_{\mu_{\rho}}$) due to the difficult position-dependent loss landscape; see Fig. 5. We keep $\lambda_{\sigma_{\rho}}$ small to avoid a collapse of the position uncertainty. For the image parameters we set $\lambda_{\mu_{\chi}}=\lambda_{\sigma_{\chi}}$ to quickly arrive at an initial image estimate, and employ a falling cosine schedule from $0.1$ to $0.001$ between steps 4,000 and 6,000 to allow the image to stabilize and to improve the effectiveness of the high-variance prior term $b_{\theta}^{\mathrm{SDE}}$ during the later optimization steps.
@@ 4.2.2 TV baseline details
A typical choice for $h_{\alpha}:[0,\infty)\to[0,\infty)$ is the Huber regularizer (see, e.g., [9, Example 4.7]),
@@ 4.2.3 RED-Diff
The algorithm we use for blind RED-Diff in this setup is given in Algorithm 2. Following [51], in the notation of Eq. 3.17 we choose the inverse SNR weighting $\tilde{\omega}(t)=\lambda_{\text{RD}}\,\frac{\sigma(t)}{\alpha(t)}.$ In [51, Prop. 2] the weight $\lambda_{\text{RD}}>0$ may also depend on the observation noise $\sigma_{\varepsilon}$. In our experiments, we choose fixed step sizes $\tau^{(l,i)}=\tau=0.1,\eta^{(l,i)}=\eta=10$ in Algorithm 2.
@@ 4.2.4 Position log-barrier
In some of the scenarios we consider, such as phase-only objects, we observe that the loss landscape for the position recovery has an unfavorable structure, which leads to positions moving off the object rather towards reasonable estimates; see Section 5.2.4. To remedy this problem, we add a 2-dimensional log-barrier loss as a hand-crafted prior on the positions, with an empirically chosen weight $\lambda_{\text{pos}}$. Namely, we consider the following barrier function, $B(s):=-\log(1-s)-\log(1+s)+\iota_{\mathbb{R}^{2}\setminus(-1,1)^{2}}(s),$ where for a set $A$, $\iota_{A}$ denotes the indicator function from convex analysis; see [58]. Our prior is then defined by an affine transformation $T$, which first shifts and scales each position $r$, i.e., $p(r)\propto\exp(-\lambda_{\text{pos}}B(T(r))$. We define $T$ such that the object extent within the domain is scaled to a smaller domain $(-a,a)^{2}\subsetneq(-1,1)^{2}$. This is done via
$T(r):=\frac{r-m}{r_{\max}-r_{\min}},$
where we set $r_{\max}=H+l,r_{\min}=H-l$ and $m=(r_{\max}+r_{\min})/2$ with $l=20$, so $l$ grants 20 pixels of leniency for the positions to lie slightly outside of the object extent. To avoid undefined gradients from the log-barrier loss, we clip the positions to lie inside the rectangle spanned by these edges before feeding them to the loss expression.
@@ 4.5 Evaluation
where $x$ is the ground-truth image and $\hat{x}$ is a reconstructed image, and we clip the estimated magnitudes to [0, 1] for evaluation to avoid large errors from isolated wrong pixels. We report aPSNR values in dB, while aSSIM takes values in [0, 1]. Since aPSNR and aSSIM ignore errors in the phase structure, we include the complex-valued metric $\operatorname{cRMS}$, introduced as $E_{o}$ in [50], which is a normalized root-mean-square error metric with a complex-valued empirical scaling factor $\gamma$ that corrects for a scale ambiguity and the global phase ambiguity inherent in phase retrieval problems:
@@ 5.1 Non-blind Baseline
We first evaluate the described algorithms in the non-blind case. While this is not the main focus of the present work, we use this simpler case as a validation of our basic methodology and to gain an impression of the reconstruction quality that could be achievable under perfect position recovery. Here we use measurement noise of variance $\sigma_{\varepsilon}^{2}=0.005$, which represents an average measurement signal-to-noise ratio (SNR) of 4.5 dB over all test objects.
In Table 1, we show the metric values of the compared methods. We can see that the variants without image priors only reach PSNR values around 12 dB, which can already be improved to around 25 dB by using a Huber-TV prior. Our variational method using the surrogate score-based prior (SSP) achieves a PSNR of 30 dB and SSIM of 0.90, closely followed by RED-Diff. This demonstrates the usefulness of the score-based data-driven priors for image quality in our problem. The optimization-based methods with no prior and a Huber-TV prior perform similarly to their VI counterparts here. For the weight parameter $\lambda$ in the Huber-TV prior, we compare $\lambda\in\{5,10,20\}$, and find that larger $\lambda$ achieve better aSSIM, though aPSNR already decreases again after $\lambda=10$. Since we aim to reconstruct complex-valued images, we use $\lambda=5$ for the VI method in all subsequent experiments, which achieves the best mean cRMS at the lowest standard deviation.
To illustrate the qualitative differences between the methods, in Fig. 3(a) we show the reconstruction with the best PSNR for each variational method. One can see that the reconstruction with no prior is heavily affected by the measurement noise, and the Huber-TV prior at $\lambda=5$ only ameliorates this to some extent. H-TV with $\lambda=10$ removes more noise but the image edges noticeably begin to blur. For a direct comparison, we show images with all three Huber-TV weights in Fig. 11. While RED-Diff seems to produce more fine details than SSP, these may be partly hallucinated, which is reflected in the slightly lower metric scores in Table 1. We note here that RED-Diff involves a weighting hyperparameter $\lambda_{\text{RD}}$ for the score-based prior loss, the choice of which is ad-hoc and affects the image in a tradeoff between adherence to the measurement and adherence to the prior. RED-Diff therefore does not sample from the actual posterior induced by the likelihood and the prior, similar to issues of the related method DPS [13] as discussed in [25].

### v2

@@ 3.2.4 Variational Minimization as Baseline
3.3 On the convergence of Algorithm 1
We briefly turn towards the convergence analysis of the scheme in Algorithm 1. This section is to be understood as a discussion of the challenges one faces in our particular setting. An in-depth study is left for future work. Moreover, we refer to [56, 41, 78] for works on the convergence analysis of other algorithms for ptychography. From optimization literature, (3.19) is known as an alternating minimization scheme (AM), block coordinate descent method or Gauss–Seidel iteration, which has been studied for example in [10, 11, 102]. We state the following standard assumption, which ensures well-definedness of the scheme:
Assumption 3.1.
The function $\mathcal{L}$ is proper, continuous and has compact sub-level sets $\{(\chi,\rho):\mathcal{L}(\chi,\rho)\leq\alpha\}$.
By definition we have that $\mathcal{L}(\chi_{k+1},\rho_{k+1})\leq\mathcal{L}(\chi_{k},\rho_{k})$ and the compactness of the sub-level sets then implies that the sequence $(\chi_{k},\rho_{k})$ is bounded. A further assumption used in [102] to ensure convergence towards stationary points is that $\mathcal{L}$ admits coordinate-wise unique minimizers.
Assumption 3.2.
[…]
Under this assumption [102, Thm. 4.1] yields the following:
Lemma 3.3.
Under Assumptions 3.2 and 3.1, any cluster point $(\chi^{*},\rho^{*})$ of $\{(\chi_{k},\rho_{k})\}$ satisfies
$\chi^{*}=\operatorname*{arg\,min}_{\chi}\mathcal{L}(\chi,\rho^{*}),\hskip 18.49988pt\rho^{*}=\operatorname*{arg\,min}_{\rho}\mathcal{L}(\chi^{*},\rho).$
If $\mathcal{L}\in C^{1}$ this means $(\chi^{*},\rho^{*})$ is a stationary point, i.e.,
$\nabla_{(\chi,\rho)}\mathcal{L}(\chi^{*},\rho^{*})=0.$
Regarding the uniqueness in the position-component, we will observe numerically in Fig. 5 that with a randomized probe one can hope to obtain this property. However, formulating the necessary assumptions to prove this statement is not trivial and outside the scope of this work. Regarding uniqueness in the object component of $\mathcal{L}$ we remark that the same ambiguities as discussed in Remark 2.1 appear. Therefore, in general, one can only hope that additional prior terms help to ensure Assumption 3.2.
In the absence of this property, the AM in Algorithm 1 is not guaranteed to converge to a stationary point and counter examples exist, see e.g. [11, Ex. 14.4]. Nevertheless, we can exploit the special structure of $\mathcal{L}$, by splitting it in the parts that depend on both $\chi$ and $\rho$ and the respectively independent parts,
$\underbrace{\mathbb{E}_{x\sim q_{\chi}}\mathbb{E}_{\mathbf{r}\sim q_{\rho}}\left[-\log p(y\,|\,x,\mathbf{r})\right]}_{=:\mathcal{F}_{1}(\chi,\rho)}+\underbrace{\mathbb{E}_{x\sim q_{\chi}}[-\log{\hat{p}_{\theta}}(x)]-\mathcal{H}(q_{\chi})}_{=:\mathcal{F}_{2}(\chi)}+\underbrace{\mathbb{E}_{\mathbf{r}\sim q_{\rho}}[-\log p(\mathbf{r})]-\mathcal{H}(q_{\rho})}_{=:\mathcal{F}_{3}(\rho)}.$
In this situation the arguments in [10] allow us to derive a convergence statement under the following assumption.
Assumption 3.4.
The functions $\mathcal{F}_{2},\mathcal{F}_{3}$ are proper, convex functions with closed sub-level sets.
We first observe that in the non-parametric setting for any positive distribution $\tilde{p}$ the function
$q\mapsto\mathbb{E}_{x\sim q}[-\log\tilde{p}(x)]-\mathcal{H}(q)$
is convex in $q$ since the first part is linear in $q$ and the entropy $\mathcal{H}$ is concave. However, when we choose a concrete parametrization, this convexity typically does not transfer to the convexity in $\chi,\rho$. Namely, in 4.5, we choose a Gaussian parametrization as follows:
$\begin{aligned} q_{\chi}(x|y):=\mathcal{N}_{\mathbb{C}}(x;\mu_{\chi},\Sigma_{\chi}I),&\ \mu_{\chi}\in\mathbb{C}^{d},&\ \Sigma_{\chi}=\operatorname{diag}(\sigma_{\chi}^{2}),&\ \sigma_{\chi}\in\mathbb{R}_{+}^{d}\,,\\ q_{\rho}(r|y):=\mathcal{N}(r;\mu_{\rho},\Sigma_{\rho}I),&\ \mu_{\rho}\in\mathbb{R}^{K\times 2},&\ \Sigma_{\rho}=\operatorname{diag}(\sigma_{\rho}^{2}),&\ \sigma_{\rho}\in\mathbb{R}_{+}^{K}\,.\end{aligned}$
The following lemma shows that this choice preserves convexity of the entropy functional.
Lemma 3.5.
For the parametrization in 4.5, the functionals $\chi\mapsto\mathcal{H}(q_{\chi})$ and $\rho\mapsto\mathcal{H}(\rho)$ are convex.
Proof.
The entropy of the Gaussian density $q_{\chi}$ is given by
$\mathcal{H}(\chi)=C+\frac{1}{2}\log\det(\operatorname{diag}(\sigma_{\chi}^{2}))=C+\frac{1}{2}\sum_{i=1}^{d}\log\,(\sigma_{\chi})_{i}^{2}$
for a constant $C$ independent of $\chi$. In particular $\mathcal{H}(\chi)$ is constant in $\mu_{\chi}$ and concave in $\sigma_{\chi}$. Thus $\chi\mapsto\mathcal{H}(q_{\chi})$ is convex, the convexity of $\rho\mapsto\mathcal{H}(\rho)$ follows analogously. ∎
In general we cannot expect convexity of the first term in $\mathcal{F}_{2},\mathcal{F}_{3}$. In order to establish this property, we would need to assume that both $x\mapsto\hat{p}_{\theta}(x)$ and $\mathbf{r}\mapsto p(\mathbf{r})$ are log-concave. We summarize this in the following lemma, which is a direct consequence from [18].
Lemma 3.6.
Assume that both $x\mapsto\hat{p}_{\theta}(x)$ and $\mathbf{r}\mapsto p(\mathbf{r})$ are log-concave functions, then $\mathcal{F}_{2},\mathcal{F}_{3}$ are convex.
Proof.
This fact is established in [18, Sec. 3]. ∎
This raises the question whether log-concavity of the priors can be assumed in our case. We first consider the position prior.
Lemma 3.7.
The function $\mathbf{r}\mapsto p(\mathbf{r})$ defined as in Section 4.2.4, i.e., the log-barrier, is log-concave.
Proof.
We have that $\log p(\mathbf{r})=\lambda_{\text{pos}}B(T(\mathbf{r}))$, where $T$ is affine and the logarithmic barrier $B$ is convex. Thus $\mathbf{r}\mapsto-\log p(\mathbf{r})$ is convex as the composition of an affine and a convex function. ∎
The data-driven image prior $\hat{p}_{\theta}$ is usually complex and multimodal, hence convexity of $\mathcal{F}_{2}$ cannot be guaranteed. A simpler setting, in which the assumptions of Lemma 3.6 are met, is reached when we replace $\hat{p}_{\theta}$ by a classical model-based prior. An example of a log-concave prior is the TV baseline in Section 4.2.2. In such a setting, we obtain the following result from [10, Lem. 3.3].
Lemma 3.8.
Assume that $\mathcal{F}_{1}\in C^{1}$ with Lipschitz continuous gradients and that both $x\mapsto\hat{p}_{\theta}(x)$ and $\mathbf{r}\mapsto p(\mathbf{r})$ are log-concave. Then, any accumulation point $(\chi^{*},\rho^{*})$ of $\{(\chi_{k},\rho_{k})\}$ is a stationary point of $\mathcal{L}$.
Note that a stationary point of $\mathcal{L}$ does not have to coincide with its minimizer. For this, $\mathcal{F}_{1}$ would need to be convex, which seems unrealistic considering Remark 2.1. While the previous notes offer only limited insights into the convergence behavior, they still highlight the main difficulties connected to the analysis. A more in-depth study of the convergence of the gradient-based schemes under priors that are not log-concave is left for future work.
@@ 4 Experiments
In the following, we describe the computational experiments we perform to arrive at our results and conclusions.
@@ 4.1 Problem Setup
First, we describe our simulated problem setup including all relevant parameters for the algorithms and the simulated measurement models. In all experiments, we use an object array of size $x\in\mathbb{C}^{H\times W}:=\mathbb{C}^{256\times 256}$ and a probe array of size $p\in\mathbb{C}^{H_{p}\times W_{p}}:=\mathbb{C}^{512\times 512}$.
@@ 4.1.4 Simulated Measurements
where each pixel in $y_{k}$ is treated as both the mean and the variance, and the square root is taken element-wise. An alternative would be to use $\left|\mathcal{A}(r_{k},x)\right|^{2}$ as the variance of the approximating Gaussian instead of $y_{k}$, but in preliminary experiments we found this option to lead to less stable reconstructions and hence always use 4.4. In practice, within the expression $\frac{1}{2y_{k}}$ we clamp $y_{k}$ to have a minimum value of 1 to avoid both division by zero and an overly high weighting of detector pixels with no measured photons.
@@ 4.2 Algorithmic Configuration
In the following, we list all relevant details for the configuration of the tested reconstruction algorithms. A summary of algorithmic parameters is provided in Table 1. Note that, due to long reconstruction runtimes, limited computational resources, and a large set of parameters to tune, we did not systematically grid-search any algorithm’s set of parameters. Instead, we performed small manual line-searches of parameters using 1-3 example images until successful reconstructions were achieved. It is not our goal here to establish any particular reconstruction method as a “state-of-the-art”, but rather to show that decent reconstructions can be attained by each method at all in this difficult problem setting, and to highlight different behaviors of the methods. Therefore, detailed tuning is out of scope for this work, and it is likely that future systematic algorithm parameter optimizations will lead to further improved reconstruction quality.
Method #Iterations $N_{\mathrm{img}}$ $N_{\mathrm{pos}}$ $\lambda_{\mathrm{img}}$ $\lambda_{\mathrm{pos}}$ $\lambda_{\sigma_{\chi}}$ $\lambda_{\sigma_{\rho}}$ $\lambda_{\mathrm{HTV}}$ $\lambda_{\text{RD}}$
VI (SSP) 10,000 1 10 0.1 10 0.1 0.01 - -
VI (H-TV prior)    5 -
VI (no prior)    - -
RED-Diff [77]    - - - 20
Opt (H-TV prior)    - - 0.1 -
Opt (No prior)    - - - -
Table 1: Algorithmic hyperparameters used for each reconstruction method in this work, unless otherwise noted. Entries with – ” – indicate that the same parameter setting is used as in the row above, and entries with - indicate that the parameter is not applicable or unused for the respective method. For easier comparison, $\lambda_{\mathrm{img}},\lambda_{\mathrm{pos}}$ respectively refer to the step sizes for the image and positions for RED-Diff and optimization-based methods, while for variational approaches they refer to the step sizes for the corresponding distribution’s mean, i.e., $\lambda_{\mu_{\chi}}$ and $\lambda_{\mu_{\rho}}$, respectively. Note that, for the VI methods, $\lambda_{\mathrm{img}}$ and $\lambda_{\sigma_{\chi}}$ are scheduled and decay starting from the listed values, see Section 4.2.1 for full details.
@@ 4.2.1 Variational Inference
For our variational inference methods and all choices of prior distribution, we approximate the posterior by the following variational distributions, which we recall from Section 3.3
$\begin{aligned} q_{\chi}(x|y):=\mathcal{N}_{\mathbb{C}}(x;\mu_{\chi},\Sigma_{\chi}I),&\ \mu_{\chi}\in\mathbb{C}^{d},&\ \Sigma_{\chi}=\operatorname{diag}(\sigma_{\chi}^{2}),&\ \sigma_{\chi}\in\mathbb{R}_{+}^{d}\,,\\ q_{\rho}(r|y):=\mathcal{N}(r;\mu_{\rho},\Sigma_{\rho}I),&\ \mu_{\rho}\in\mathbb{R}^{K\times 2},&\ \Sigma_{\rho}=\operatorname{diag}(\sigma_{\rho}^{2}),&\ \sigma_{\rho}\in\mathbb{R}_{+}^{K}\,,\\ q_{\chi,\rho}(x,r|y):=q_{\chi}(x|y)q_{\rho}(r|y)\,.\end{aligned}$ (4.5)
@@
For both the image and position parameters, we use the Adam optimizer [65], which has previously been established as a working method for (non-blind) ptychographic image reconstruction [61, 59, 54]. We set its momentum parameters to $\beta_{1}=0.9,\beta_{2}=0.999$, the default settings in the PyTorch package [83]. We run $N=\text{10,000}$ outer steps unless otherwise noted, each with $N_{\mathrm{img}}=1$ inner image optimization step and $N_{\mathrm{par}}=10$ inner position optimization steps. For the position parameters we set the step sizes $\lambda_{\mu_{\rho}}=10.0$ and $\lambda_{\sigma_{\rho}}=0.01$, where we use a high learning rate of the means ($\lambda_{\mu_{\rho}}$) due to the difficult position-dependent loss landscape; see Fig. 5. We keep $\lambda_{\sigma_{\rho}}$ small to avoid a collapse of the position uncertainty. For the image parameters we set $\lambda_{\mu_{\chi}}=\lambda_{\sigma_{\chi}}$ to quickly arrive at an initial image estimate, and employ a falling cosine schedule from $0.1$ to $0.001$ between steps 4,000 and 6,000 to allow the image to stabilize and to improve the effectiveness of the high-variance prior term $b_{\theta}^{\mathrm{SDE}}$ during the later optimization steps.
@@ 4.2.2 TV baseline details
A typical choice for $h_{\alpha}:[0,\infty)\to[0,\infty)$ is the Huber regularizer (see, e.g., [19, Example 4.7]),
@@ 4.2.3 RED-Diff
The algorithm we use for blind RED-Diff in this setup is given in Algorithm 2. Following [77], in the notation of Eq. 3.21 we choose the inverse SNR weighting $\tilde{\omega}(t)=\lambda_{\text{RD}}\,\frac{\sigma(t)}{\alpha(t)}.$ In [77, Prop. 2] the weight $\lambda_{\text{RD}}>0$ may also depend on the observation noise $\sigma_{\varepsilon}$. In our experiments, we choose fixed step sizes $\tau^{(l,i)}=\tau=0.1,\eta^{(l,i)}=\eta=10$ in Algorithm 2 and $\lambda_{\text{RD}}=20$.
@@ 4.2.4 Position log-barrier
In some of the scenarios we consider, such as phase-only objects, we observe that the loss landscape for the position recovery has an unfavorable structure, which leads to positions moving off the object rather towards reasonable estimates; see Section 5.2.5. To remedy this problem, we add a 2-dimensional log-barrier loss as a hand-crafted prior on the positions, with an empirically chosen weight $\lambda_{\text{pos}}$. Namely, we consider the following barrier function, $B(s):=-\log(1-s)-\log(1+s)+\iota_{\mathbb{R}^{2}\setminus(-1,1)^{2}}(s),$ where for a set $A$, $\iota_{A}$ denotes the indicator function from convex analysis; see [86]. Our prior is then defined by an affine transformation $T$, which first shifts and scales each position $r$, i.e., $p(r)\propto\exp(-\lambda_{\text{pos}}B(T(r))$. We define $T$ such that the object extent within the domain is scaled to a smaller domain $(-a,a)^{2}\subsetneq(-1,1)^{2}$. This is done via $T(r):=\frac{r-m}{r_{\max}-r_{\min}}$, where we set $r_{\max}=H+l,r_{\min}=H-l$ and $m=(r_{\max}+r_{\min})/2$ with $l=20$, so $l$ grants 20 pixels of leniency for the positions to lie slightly outside of the object extent. To avoid undefined gradients from the log-barrier loss, we clip the positions to lie inside the rectangle spanned by these edges before feeding them to the loss expression.
@@ 4.5 Evaluation
where $x$ is the ground-truth image and $\hat{x}$ is a reconstructed image, and we clip the estimated magnitudes to [0, 1] for evaluation to avoid large errors from isolated wrong pixels. For all methods based on variational inference, we treat the final fitted distributional mean $\mu_{\chi}$ as the image estimate $\hat{x}$ for evaluation, and for producing the images shown in the figures in Section 5. We report aPSNR values in dB, while aSSIM takes values in [0, 1]. Since aPSNR and aSSIM ignore errors in the phase structure, we include the complex-valued metric $\operatorname{cRMS}$, introduced as $E_{o}$ in [76], which is a normalized root-mean-square error metric with a complex-valued empirical scaling factor $\gamma$ that corrects for a scale ambiguity and the global phase ambiguity inherent in phase retrieval problems:
@@ 5.1 Non-blind Baseline
We first evaluate the described algorithms in the non-blind case. While this is not the main focus of the present work, we use this simpler case as a validation of our basic methodology and to gain an impression of the reconstruction quality that could be achievable under perfect position recovery. Here we use measurement noise of standard deviation $\sigma_{\varepsilon}=0.005$, which represents an average measurement signal-to-noise ratio (SNR) of 4.5 dB over all test objects.
In Table 2, we show the metric values of the compared methods. We can see that the variants of VI and optimization-based methods without image priors only reach PSNR values around 12 dB, which can already be improved to around 25 dB by using a Huber-TV prior. The variational surrogate score-based prior (SSP) method achieves a PSNR of 30 dB and SSIM of 0.90, closely followed by RED-Diff. This demonstrates the usefulness of the score-based data-driven priors for image quality in our problem. The optimization-based methods with no prior and a Huber-TV prior perform similarly to their VI counterparts here. For the weight parameter $\lambda$ in the Huber-TV prior, we compare $\lambda\in\{5,10,20\}$, and find that larger $\lambda$ achieve better aSSIM, though aPSNR already decreases again after $\lambda=10$. Since we aim to reconstruct complex-valued images, we use $\lambda=5$ for the VI method in all subsequent experiments, which achieves the best mean cRMS at the lowest standard deviation.
To illustrate the qualitative differences between the methods, in Fig. 3(a) we show the reconstruction with the best PSNR for each variational method. One can see that the reconstruction with no prior is heavily affected by the measurement noise, and the Huber-TV prior at $\lambda=5$ only ameliorates this to some extent. As shown in Fig. 14, H-TV with $\lambda=10$ removes more noise but the image edges noticeably begin to blur. While RED-Diff seems to produce more fine details than SSP, these may be partly hallucinated, which is reflected in the slightly lower metric scores in Table 2. We note here that RED-Diff involves a weighting hyperparameter $\lambda_{\text{RD}}$ for the score-based prior loss, the choice of which is ad-hoc and affects the image in a tradeoff between adherence to the measurement and adherence to the prior. RED-Diff therefore does not sample from the actual posterior induced by the likelihood and the prior, similar to issues of the related method DPS [25] as discussed in [39].